A fighter jet passes directly above your head. A few moments later, when the imaginary line from your eye to the jet makes a 30 degree angle with the ground you hear a sonic boom. What is the speed of the aircraft?
Answer #1

The half angle of Mach cone is

\sin\theta=v/vs

where v is speed of sound and vs is speed of aircraft.

so=vs= v/ \sin\theta=343/\sin30=686 m/s